OpenStudy (anonymous):

Find the distance traveled by an object on the interval [0, 5] if its velocity is v(t)= sin ( pie/3 t) a. 21/2pie b. 3/2pie c. 7/2 d. 1/2 It's multiple choice of the above from a - d which one is the answer? I know it's intergrals, and I have to compute the intergral of the function and plug in the points on the intergral. But my calculator was refusing the compute Help anyone?

OpenStudy (anonymous):

you are doing math with food i think its spelt pi

OpenStudy (anonymous):

LOL sorry! yes it's pi I was in a hurry. My bad.

OpenStudy (lgbasallote):

sounds like an integral questin

OpenStudy (anonymous):

shold be b

OpenStudy (lgbasallote):

\[\huge \int_0^5 \sin (\frac{\pi}{3t}) dt\] is that how it looks like?

OpenStudy (lgbasallote):

or \[\huge\int_0^5 \sin (\frac{\pi}{3} t) dt\]

OpenStudy (anonymous):

This is how it looks like.

OpenStudy (anonymous):

\[\int_0^3\sin(\frac{\pi t}{3})dt-\int_3^5\sin(\frac{\pi t}{3})dt\]

OpenStudy (anonymous):

reason for breaking up the interval in to two parts is that first is is going forward and then it is going backwards

OpenStudy (anonymous):

Thanks man! :D you are one smart one! :D means a lot!